OpenStudy (anonymous):

The position s of a point (in feet) is given as a function of time t (in seconds). s = 2/t − 2/t^2; t = 2 (a) Find its acceleration as a function of t. s''(t) = (b) Find the point's acceleration at the specified time. s''(2) =

OpenStudy (anonymous):

$$a(x) = \frac{d^2 s}{dt^2}$$

OpenStudy (anonymous):

Its s= \[\frac{ 2 }{ t }-\frac{ 2 }{ t^2 }\] t=2

OpenStudy (anonymous):

s''= \[(4/t^{3})-(12/t ^{4})\] s''(2)= -0.25 m/s^2

OpenStudy (anonymous):

@archanatikayatray its actually 4/t^3-12/t^4 Thank you!

OpenStudy (anonymous):

@arshia93 You're Welcome